OpenStudy (anonymous):

solve the equation by completing the square a^2+2a-35=0

OpenStudy (anonymous):

a^2+2a-35=0 (a+7)(a-5)=0 (By factorization) take (=0) for both sides, a=-7 or a=5

OpenStudy (amistre64):

Completing the square: a^2 +2a -35=0 -- add (35); a^2 +2a +K = 35 +K -- K = (b/2)^2 where b=2 a^2 +2a +1 = 35+1 -- factor the LHS into a perfect square; add RHS together (a+1)^2 = 36 -- take the square root of both sides a+1 = 6 or a+1 = -6 --solve for "a" a= 5 or a=(-7) -- check your answers to eliminate false hope :) (5)^2 +2(5) -35 = 0 25 +10 -35 = 0 35 - 35 = 0 (a=5, is good) (-7)^2 +2(-7) -35 = 0 49 - 14 - 35 = 0 49 - 49 = 0 (a=(-7), is good) a= -7 and a=5

OpenStudy (amistre64):

On a test, I would just factor like razmik did to get the answers; then double check :)